# The Achilles’ Heel of LLMs: How Altering a Handful of Neurons Can Cripple Language Abilities

### v1

@@ Abstract
Large Language Models (LLMs) have become foundational tools in natural language processing, powering a wide range of applications and research. Many studies have shown that LLMs share significant similarities with the human brain. Recent neuroscience research has found that a small subset of biological neurons in the human brain are crucial for core cognitive functions, which raises a fundamental question: do LLMs also contain a small subset of critical neurons? In this paper, we investigate this question by proposing a Perturbation-based Causal Identification of Critical Neurons method to systematically locate such critical neurons in LLMs. Our findings reveal three key insights: (1) LLMs contain ultra-sparse critical neuron sets. Disrupting these critical neurons can cause a 72B-parameter model with over 1.1 billion neurons to completely collapse, with perplexity increasing by up to 20 orders of magnitude; (2) These critical neurons are not uniformly distributed, but tend to concentrate in the outer layers, particularly within the MLP down_proj components; (3) Performance degradation exhibits sharp phase transitions, rather than a gradual decline, when these critical neurons are disrupted. Through comprehensive experiments across diverse model architectures and scales, we provide deeper analysis of these phenomena and their implications for LLM robustness and interpretability. These findings can offer guidance for developing more robust model architectures and improving deployment security in safety-critical applications.
@@ 1 Introduction
Understanding the fundamental principles underlying intelligent computation has long attracted researchers across neuroscience and artificial intelligence. A growing body of evidence suggests that the brain’s computational power may critically depend on a small fraction of highly influential neurons, rather than being uniformly distributed across all neural elements. This principle, first stated by  Arshavsky (2001), suggests that sparsely distributed, genetically distinct neurons function as computational bottlenecks, each having a major influence over specific brain functions.
Recent studies have provided support for this hypothesis across multiple neural systems (Meissner-Bernard et al., 2025; Fan et al., 2025). In the hippocampus, researchers have found small groups of neurons that work together during specific brain events. The neuron groups show precise timing when they activate together. When disrupting these critical neuron groups, memory formation becomes severely impaired, though these groups represent only a tiny fraction of all neurons in the brain (Kunz et al., 2024). Similarly, in sensory cortex, sparse neurons using predictive coding mechanisms prove essential for efficient perception, with their dysfunction causing pathological states including epileptic activity (Chalk et al., 2018). Motor systems show the same pattern of dependency on sparse neurons. Brief activation patterns in small neuron groups control complex movement sequences. When researchers silence these groups, immediate motor problems occur (Zhang et al., 2024). These findings together suggest that biological neural networks operate under a “sparse criticality” principle, where cognitive and behavioral functions depend critically on the coordinated activity of a minority of highly influential neurons.
In parallel, Large Language Models (LLMs)—with their billions of parameters—have demonstrated capabilities mirroring human cognition, from language comprehension to problem-solving (Riztha et al., 2024; Niu et al., 2024). Recent research shows clear parallels: the hierarchical processing and predictive coding mechanisms in LLMs align closely with the brain’s layered neural architectures (Schrimpf et al., 2021; Caucheteux & King, 2022; Mischler et al., 2024), hinting that these artificial systems may emulate biological computation. This observation leads to a key question:
@@
Figure 1: Illustration of critical neuron identification and progressive masking effects, using DeepSeek-R1-Distill-Llama-70B as an example. The top panel shows that our method identifies 3 critical neurons located in MLP down_proj components of the transformer architecture. The bottom panel demonstrates the progressive degradation of model performance on the question "What is the capital of France?" through sequential masking: the left chart shows token probabilities after masking the first critical neuron, the middle chart shows the effect of masking the first two critical neurons, and the right chart reveals catastrophic failure after masking all three critical neurons. The progression illustrates a sharp phase transition where masking the complete critical neuron set triggers sudden collapse rather than gradual degradation.
In neuroscience, the causal necessity of specific neurons for a given function is often established through lesion studies, which assess behavioral changes following the removal or deactivation of targeted brain regions (Vaidya et al., 2019). Inspired by this classic experimental approach, we ask whether LLMs also show a comparable “critical neuron" phenomenon. To explore this possibility, we similarly adapt the logic of lesion studies to LLMs: since the core function of an LLM is to transform input text into meaningful output for natural language processing tasks (Qin et al., 2024; Zhao et al., 2025), we need to examine the quality and coherence of model outputs after deactivating specific neurons.
@@
To put this investigation into practice, we introduce the Perturbation-based Causal Identification of Critical Neurons method, as illustrated in Figure 1. Given any text as input, the proposed method consists of a two stage optimization process that mirrors the spirit of lesion studies to locate the so called critical neurons. In the first stage, the method injects controlled noise into the model’s input and measures the resulting activation differences across neurons, thereby generating a ranked list of candidates most likely to influence model behavior. In the second stage, we sequentially mask these top-ranked neurons in a greedy manner, closely monitoring changes in model perplexity. This allows us to causally determine which neurons are truly essential for the model’s function.
@@
Using this method, we analyze a diverse set of datasets—including Neurons, WikiText-103, C4, MMLU-Pro (Wang et al., 2025), IFEval (Zhou et al., 2023), GPQA-Diamond (Rein et al., 2024), HumanEval (Peng et al., 2024), MATH (Manem et al., 2025), MGSM (Shi et al., 2023), and SimpleQA (Wei et al., 2024)—as well as multiple LLMs such as Llama-3 (Grattafiori et al., 2024), Gemma (Gemma Team, 2024), DeepSeek-R1 (DeepSeek-AI, 2025), Phi-3, Phi-3.5 (Abdin et al., 2024), and Qwen2.5 (Qwen, 2025). According to our experiment, we observe consistent and interesting patterns across all settings. In particular, our findings reveal several key phenomena. First, we discover that LLMs are governed by an ultra-sparse set of critical neurons: astonishingly, disabling as few as three neurons can catastrophically impair a 72B-parameter model comprising over 1.1 billion neurons, driving its perplexity up by as much as 20 orders of magnitude. Second, we find that these critical neurons are not randomly distributed throughout the network, but are instead highly concentrated in the outer layers, particularly within the MLP down_proj components, which is consistent with the observations reported in Yu et al. (2025). Third, we observe that performance degradation does not occur gradually, but rather through sharp phase transitions. In section 4, we provide a detailed discussion of these findings and hope these discoveries can help the community better optimize LLM architecture, training, and inference processes. Our contributions are summarized as follows:
@@
  • Inspired by lesion studies in neuroscience, we propose a systematic method that integrates noise-based sensitivity analysis with causal verification, which enables us precisely to identify critical neurons that are indispensable for the overall functionality of LLMs. Our method shows exceptional robustness, consistently identifying identical critical neurons for any input texts exceeding a minimum token length threshold ($T>10$).
@@ 2 Related Work
Critical Components Analysis in LLMs. Recent research has revealed a variety of vulnerabilities in LLMs, but most studies have focused on parameter-level rather than neuron-level analysis. For example, Kovaleva et al. (2021) investigates weight outliers in LLMs and finds that anomalous parameters arising during pre-training can significantly affect performance when disabled. Bondarenko et al. (2021) explores activation outliers that influence attention mechanisms, particularly in processing special tokens. Sun et al. (2024) uncovers persistent, large-scale activation patterns at fixed positions across transformer layers, and Yang et al. (2024) attributes such phenomena to architectural components like gated linear units. At the same time, studies on “super weights” show that individual parameters in MLP down_proj layers can catastrophically impact model performance (Yu et al., 2025), though this line of work mainly addresses quantization rather than systematic vulnerability analysis. While these works together reveal important architectural sensitivities and suggest the existence of critical neurons, they fall short in two key aspects: first, there is still a lack of effective methods for accurately identifying these key neurons; second, systematic analysis of their properties and roles within the model remains unexplored. Our work addresses these limitations by proposing approaches to locate critical neurons and by conducting a comprehensive analysis of their functions across diverse model architectures and scales.
@@
Neurons Localization and Editing. Recent studies have advanced neuron-level localization and editing in LLMs, showing that neural parameters can be divided between pattern learning and memorization (Bender et al., 2021). Targeted interventions on specific neurons or clusters enable systematic modification of factual knowledge (Meng et al., 2022; Dai et al., 2022; Li et al., 2024), control of personal information memorization (Chen et al., 2024), multilingual capabilities (Tang et al., 2024), and even safety constraints (Zhou et al., 2025). Further, behavioral attributes such as “aggression” can be manipulated by localizing and editing a small set of neurons (Lee et al., 2025), and polysemantic neurons in fusion heads are linked to modality collapse in multimodal models (Chaudhuri et al., 2025). However, existing work primarily focuses on task-specific interventions, lacking systematic frameworks for identifying neurons critical to core model functionality. Our work addresses this gap by introducing a principled method for neuron localization and causal verification, inspired by lesion studies in neuroscience.
@@ 4.1 Experimental Setup
We conduct our investigation across multiple datasets to assess both language modeling capabilities and downstream task performance. For language modeling, we primarily use WikiText-103 and the C4 dataset. To evaluate the impact of critical neuron masking on specific capabilities, we employ seven additional benchmarks: MMLU Pro (Wang et al., 2025) for advanced multi-domain knowledge, IFEval (Zhou et al., 2023) for instruction following, GPQA Diamond (Rein et al., 2024) for graduate-level scientific reasoning, HumanEval (Peng et al., 2024) for code generation, MATH (Manem et al., 2025) for mathematical problem solving, SimpleQA (Wei et al., 2024) for factual question answering, and MGSM (Shi et al., 2023) for multilingual mathematical reasoning. Detailed evaluation protocols and metrics for each benchmark are provided in Appendix A.3.
@@ 4.2 Investigation Results
This concentration reflects the distinct computational roles of different network depths. Early layers handle fundamental feature extraction, making them critical for establishing the computational foundation. Late layers perform final integration and output generation, serving as bottlenecks for translating representations into responses. Middle layers perform step-by-step refinement that can tolerate disruption through redundant pathways (Sun et al., 2025; Bogomasov & Conrad, 2025). MLP down_proj components in outer layers are particularly vulnerable because they compress high-dimensional representations, creating information bottlenecks. When these compression points fail, information loss spreads throughout the network, causing system-wide collapse.
@@
The sharp phase transition behavior can be attributed to the highly interdependent nature of information processing in transformer architectures. Critical neurons appear to form a tightly coupled computational circuit where each neuron’s function depends critically on inputs from others within the same set. Individual neuron masking fails to disrupt the circuit because remaining neurons can partially compensate through their shared computational pathways. When multiple critical neurons are simultaneously removed, the circuit loses its functional integrity, creating cascading information bottlenecks that spread throughout the network (Ziyin & Ueda, 2022; Manem et al., 2025).
4.3 Comparison with Existing Neuron Localization Methods
To validate the effectiveness of our method, we conduct comparison studies examining different neuron location strategies. We systematically compare our approach against alternative neuron location strategies across three representative 70B models. Our evaluation includes random selection and two importance-based alternatives: activation magnitude (AM) (Sattarifard et al., 2025) ranking and gradient magnitude (GM) (Liu et al., 2025) ranking. We evaluate each strategy by progressively masking neurons from 0 to 1,000 and measuring the perplexity. AM ranks neurons by their activation values, while GM ranks by the magnitude of perplexity gradients with respect to each neuron. For random selection, we average results over 10 trials to ensure statistical reliability.
@@ 4.4 Ablation Study
Our method demonstrates exceptional robustness. Particularly, we demonstrate that for input texts exceeding a minimum token length threshold ($T>10$), regardless of text types (Wikipedia, news, biography, media) or languages (English, French, German, Chinese, Spanish), the identified critical neurons remain consistent for each model when $K$ and $\alpha$ reach stable values ($K=100$, $\alpha=5$). Furthermore, we find that models subjected to fine-tuning and reinforcement learning retain the same critical neuron locations, and masking these neurons continues to cause identical catastrophic failures, indicating that these architectural dependencies persist across different training approaches. The token length threshold analysis, detailed cross-linguistic validation results, and fine-tuning robustness experiments are provided in Appendix A.7.
We conduct parameter analysis to examine the stability of our method. Figure 4 shows that the noise scale $\alpha$ exhibits a clear threshold effect, with fluctuations before $\alpha=5$, then stabilizing consistently at 7 neurons for all values where $\alpha\geq 5$. The sample size $K$ demonstrates different convergence characteristics, showing initial fluctuations at smaller sample sizes but achieving consistent stability at 7 neurons once $K$ exceeds 80. These convergence patterns reveal that parameters have distinct minimum thresholds: $\alpha$ requires sufficient perturbation to effectively distinguish critical neurons from non-critical ones, while $K$ needs adequate sampling to reduce statistical noise. After carefully balancing computational efficiency and accuracy requirements, we select $\alpha=5$ and $K=100$ as our standard parameters.
@@
Additionally, we validate that completely masking critical neurons represents the most effective perturbation strategy compared to alternative scaling approaches. Through analysis of different scaling factors applied to critical neuron activations, we confirm that complete masking produces the maximum performance degradation, justifying our methodological choice (see Appendix A.8).

### v2

@@ Abstract
Large Language Models (LLMs) have become foundational tools in natural language processing, powering a wide range of applications and research. Many studies have shown that LLMs share significant similarities with the human brain. Neuroscience research has found that a small subset of biological neurons in the human brain are crucial for core cognitive functions, which raises a fundamental question: do LLMs also contain a small subset of critical neurons? In this paper, we investigate this question by proposing a Perturbation-based Causal Identification of Critical Neurons method to systematically locate such critical neurons in LLMs. Our findings reveal three key insights: (1) LLMs contain ultra-sparse critical neuron sets. Disrupting these critical neurons can cause a 72B-parameter model with over 1.1 billion neurons to completely collapse, with perplexity increasing by up to 20 orders of magnitude; (2) These critical neurons are not uniformly distributed, but tend to concentrate in the outer layers, particularly within the MLP down_proj components; (3) Performance degradation exhibits sharp phase transitions, rather than a gradual decline, when these critical neurons are disrupted. Through comprehensive experiments across diverse model architectures and scales, we provide deeper analysis of these phenomena and their implications for LLM robustness and interpretability. These findings can offer guidance for developing more robust model architectures and improving deployment security in safety-critical applications. Our code is available at https://github.com/qqqqqqqzx/The-Achilles-Heel-of-LLMs.
@@ 1 Introduction
Understanding the fundamental principles underlying intelligent computation has long attracted researchers across neuroscience and artificial intelligence. A growing body of evidence suggests that the brain’s computational power may critically depend on a small fraction of highly influential neurons, rather than being uniformly distributed across all neural elements. This principle, stated by  Arshavsky (2001), suggests that sparsely distributed, genetically distinct neurons function as computational bottlenecks, each having a major influence over specific brain functions.
In parallel, Large Language Models (LLMs)—with their billions of parameters—have demonstrated capabilities mirroring human cognition, from language comprehension to problem-solving (Riztha et al., 2024; Niu et al., 2024). Recent research shows clear parallels: the hierarchical processing and predictive coding mechanisms in LLMs align closely with the brain’s layered neural architectures (Schrimpf et al., 2021; Caucheteux and King, 2022; Mischler et al., 2024), hinting that these artificial systems may emulate biological computation. This observation leads to a key question:
@@
In the field of artificial intelligence, previous work has used perturbation-based methods to identify important model components (Hanna et al., 2023; Zhang and Nanda, 2024). To explore this key question, we apply perturbation-based analysis to identify critical neurons. Our goal is to find a small subset of neurons whose collective disruption causes catastrophic system-wide collapse.
@@
To put this investigation into practice, we introduce the Perturbation-based Causal Identification of Critical Neurons method, as illustrated in Figure 1. Given any text as input, the proposed method consists of a two stage optimization process to locate the so called critical neurons. In the first stage, the method injects controlled noise into the model’s input and measures the resulting activation differences across neurons, thereby generating a ranked list of candidates most likely to influence model behavior. In the second stage, we sequentially mask these top-ranked neurons in a greedy manner, closely monitoring changes in model perplexity. This allows us to causally determine which neurons are truly essential for the model’s function.
Figure 1: Illustration of critical neuron identification and progressive masking effects, using DeepSeek-R1-Distill-Llama-70B as an example. The top panel shows that our method identifies 3 critical neurons located in MLP down_proj components of the transformer architecture. The bottom panel demonstrates the progressive degradation of model performance on the question "What is the capital of France?" through sequential masking: the left chart shows token probabilities after masking the first critical neuron (Paris probability: 0.737793, rank 1), the middle chart shows the effect of masking the first two critical neurons (Paris probability: 0.457275, rank 1), and the right chart reveals catastrophic failure after masking all three critical neurons (Paris probability: 0.000005, rank 30698). The progression illustrates a sharp phase transition where masking the complete critical neuron set triggers sudden collapse rather than gradual degradation.
@@
Using this method, we analyze a diverse set of datasets—including Neurons, WikiText-103, C4, MMLU-Pro (Wang et al., 2025), IFEval (Zhou et al., 2023), GPQA-Diamond (Rein et al., 2024), HumanEval (Peng et al., 2024), MATH (Hendrycks et al., 2021), MGSM (Shi et al., 2023), and SimpleQA (Wei et al., 2024)—as well as multiple LLMs such as Llama-3 (Grattafiori et al., 2024), Gemma (Gemma Team, 2024), DeepSeek-R1 (DeepSeek-AI, 2025), Phi-3, Phi-3.5 (Abdin et al., 2024), and Qwen2.5 (Qwen, 2025). According to our experiment, we observe consistent and interesting patterns across all settings. In particular, our findings reveal several key phenomena. First, we discover that LLMs are governed by an ultra-sparse set of critical neurons: astonishingly, disabling as few as three neurons can catastrophically impair a 72B-parameter model comprising over 1.1 billion neurons, driving its perplexity up by as much as 20 orders of magnitude. Second, we find that these critical neurons are not randomly distributed throughout the network, but are instead highly concentrated in the outer layers, particularly within the MLP down_proj components, which is consistent with the observations reported in Yu et al. (2025). Third, we observe that performance degradation does not occur gradually, but rather through sharp phase transitions. In section 4, we provide a detailed discussion of these findings and hope these discoveries can help the community better optimize LLM architecture, training, and inference processes. Our contributions are summarized as follows:
@@
  • We propose a systematic method that integrates noise-based sensitivity analysis with causal verification, which enables us precisely to identify critical neurons that are indispensable for the overall functionality of LLMs. Our method shows exceptional robustness, consistently identifying identical critical neurons for any input texts exceeding a minimum token length threshold ($T>10$).
@@ 2 Related Work
Critical Components Analysis in LLMs. Recent research has revealed a variety of vulnerabilities in LLMs, but most studies have focused on parameter-level or component-level rather than neuron-level analysis. For example, Kovaleva et al. (2021) investigates weight outliers in LLMs and finds that anomalous parameters arising during pre-training can significantly affect performance when disabled. Bondarenko et al. (2021) explores activation outliers that influence attention mechanisms, particularly in processing special tokens. Sun et al. (2024) uncovers persistent, large-scale activation patterns at fixed positions across transformer layers, and Yang et al. (2024) attributes such phenomena to architectural components like gated linear units. At the same time, studies on “super weights” show that individual parameters in MLP down_proj layers can catastrophically impact model performance (Yu et al., 2025). Our work goes beyond these studies by introducing a neuron-level perspective, identified the "critical neurons", and ruled out the possibility that they were scale outliers. We propose a Perturbation-based Causal Identification framework to systematically locate these critical neurons. Though our work may appear similar to that of Yu et al. (2025), it is fundamentally different in the following aspects: (1) Disruption strength: Their weight-level perturbations only reduce accuracy, whereas our neuron-level lesions cause a complete loss of language capability; (2) Methodological applicability: Their identification process requires manual observation, while our automated Perturbation-based Causal Identification framework provides a reproducible and engineering-ready pipeline; (3) Multi-neuron interaction: Their analysis does not examine interaction effects, whereas we reveal a phase-transition phenomenon triggered only by collective neuron disruption; (4) Research unit: They study parameters, whereas we analyze neurons; (5) Discovery process: Our findings emerge from analyzing functional collapse patterns, which reveal sharp phase transitions when coordinated neuron groups are disrupted.
@@
Neurons Localization and Editing. Recent studies have advanced neuron-level localization and editing in LLMs, showing that neural parameters can be divided between pattern learning and memorization (Bender et al., 2021). Targeted interventions on specific neurons or clusters enable systematic modification of factual knowledge (Meng et al., 2022; Dai et al., 2022; Li et al., 2024b), control of personal information memorization (Chen et al., 2024), multilingual capabilities (Tang et al., 2024), and even safety constraints (Zhou et al., 2025). Recent work on “super-neurons” (Gong et al., 2025) instead focuses on neurons with extreme polysemantic load, showing that amplifying their activations can strongly steer behavior, while masking them has only a modest effect. Related recent work on “Wasserstein neurons” (Kong et al., 2025) identifies neurons that encode syntactic structure via non-Gaussian negative pre-activations and demonstrates that disrupting many such neurons selectively impairs syntactic competence. Further, behavioral attributes such as “aggression” can be manipulated by localizing and editing a small set of neurons (Lee et al., 2025), and polysemantic neurons in fusion heads are linked to modality collapse in multimodal models (Chaudhuri et al., 2025). However, existing work primarily focuses on task-specific interventions, lacking systematic frameworks for identifying neurons critical to core model functionality. Our work addresses this gap by introducing a principled method for neuron localization and causal verification.
@@ 4.1 Experimental Setup
We conduct our investigation across multiple datasets to assess both language modeling capabilities and downstream task performance. For language modeling, we primarily use WikiText-103 and the C4 dataset. To evaluate the impact of critical neuron masking on specific capabilities, we employ seven additional benchmarks: MMLU Pro (Wang et al., 2025) for advanced multi-domain knowledge, IFEval (Zhou et al., 2023) for instruction following, GPQA Diamond (Rein et al., 2024) for graduate-level scientific reasoning, HumanEval (Peng et al., 2024) for code generation, MATH (Hendrycks et al., 2021) for mathematical problem solving, SimpleQA (Wei et al., 2024) for factual question answering, and MGSM (Shi et al., 2023) for multilingual mathematical reasoning. Detailed evaluation protocols and metrics for each benchmark are provided in Appendix A.3.
@@ 4.2 Investigation Results
This concentration reflects the distinct computational roles of different network depths. Early layers handle fundamental feature extraction, making them critical for establishing the computational foundation. Late layers perform final integration and output generation, serving as bottlenecks for translating representations into responses. Middle layers perform step-by-step refinement that can tolerate disruption through redundant pathways (Sun et al., 2025; Bogomasov and Conrad, 2025). MLP down_proj components in outer layers are particularly vulnerable because they compress high-dimensional representations, creating information bottlenecks. When these compression points fail, information loss spreads throughout the network, causing system-wide collapse.
@@
The sharp phase transition behavior can be attributed to the highly interdependent nature of information processing in transformer architectures. Critical neurons appear to form a tightly coupled computational circuit where each neuron’s function depends critically on inputs from others within the same set. Individual neuron masking fails to disrupt the circuit because remaining neurons can partially compensate through their shared computational pathways. When multiple critical neurons are simultaneously removed, the circuit loses its functional integrity, creating cascading information bottlenecks that spread throughout the network (Ziyin and Ueda, 2022).
4.3 Comparison Experiments
Comparison with Existing Neuron Localization Methods. To validate the effectiveness of our method, we conduct comparison studies examining different neuron location strategies. We systematically compare our approach against alternative neuron location strategies across three representative 70B models. Our evaluation includes random selection and two importance-based alternatives: activation magnitude (AM) (Sattarifard et al., 2025) ranking and gradient magnitude (GM) (Liu et al., 2025) ranking. We evaluate each strategy by progressively masking neurons from 0 to 1,000 and measuring the perplexity. AM ranks neurons by their activation values, while GM ranks by the magnitude of perplexity gradients with respect to each neuron. For random selection, we average results over 10 trials to ensure statistical reliability.
@@ 4.3 Comparison Experiments
Comparison with Other Components. To further clarify where the catastrophic effect originates, we compare the sensitivity of different MLP components and different transformer components.
As shown in Table  3, masking neurons in the up_proj and gate_proj according to their $\text{Imp}(s)$ causes only mild degradation, with perplexity increases remaining within 0–1 orders of magnitude even when hundreds or thousands neurons are masked.
Table 3: Perplexity before and after masking the top 10, top 100, and top 1000 most important neurons in the early (up_proj) and middle (gate_proj) MLP layers. "Baseline" is the original perplexity without masking; each column shows the average perplexity on 1,000 WikiText-103 samples after masking the specified number of neurons ranked by $\text{Imp}(s)$.
Model Baseline PPL Mask Top-10 Mask Top-100 Mask Top-1000
Qwen2.5-0.5B-Instruct 18.1606 40.2923 40.1173 43.1215
Llama-3.2-1B-Instruct 17.9204 73.5823 77.6493 307.3553
Llama-3.2-3B-Instruct 13.4696 16.5175 16.5206 497.5569
In contrast, masking only a few neurons in the MLP down_proj leads to a collapse of up to 5-6 orders of magnitude. This clear contrast shows that catastrophic failure is highly localized to the MLP down_proj component rather than a general property of transformer or MLP neurons. The need for many more neurons in other components to induce comparable collapse strengthens the causal evidence for MLP down_proj neurons (further details and analysis are provided in the Appendix A.9).
@@ 4.4 Ablation Study
Parameter Sensitivity Analysis. Our method demonstrates exceptional robustness. Particularly, we demonstrate that for input texts exceeding a minimum token length threshold ($T>10$), regardless of text types (Wikipedia, news, biography, media) or languages (English, French, German, Chinese, Spanish), the identified critical neurons remain consistent for each model when $K$ and $\alpha$ reach stable values ($K=100$, $\alpha=5$). Furthermore, we find that models subjected to fine-tuning and reinforcement learning retain the same critical neuron locations, and masking these neurons continues to cause identical catastrophic failures, indicating that these architectural dependencies persist across different training approaches. The token length threshold analysis, detailed cross-linguistic validation results, and fine-tuning robustness experiments are provided in Appendix A.7.
@@
Figure 4 shows that the noise scale $\alpha$ exhibits a clear threshold effect, with fluctuations before $\alpha=5$, then stabilizing consistently at 7 neurons for all values where $\alpha\geq 5$. The sample size $K$ demonstrates different convergence characteristics, showing initial fluctuations at smaller sample sizes but achieving consistent stability at 7 neurons once $K$ exceeds 80. These convergence patterns reveal that parameters have distinct minimum thresholds: $\alpha$ requires sufficient perturbation to effectively distinguish critical neurons from non-critical ones, while $K$ needs adequate sampling to reduce statistical noise. After carefully balancing computational efficiency and accuracy requirements, we select $\alpha=5$ and $K=100$ as our parameters.
Effectiveness of Complete Masking. We validate that completely masking critical neurons represents the most effective perturbation strategy compared to alternative scaling approaches. Through analysis of different scaling factors applied to critical neuron activations, we confirm that complete masking produces the maximum performance degradation, justifying our methodological choice (see Appendix A.8).
Precision Robustness. To rule out the possibility that the identified critical neurons are merely scale outliers, we conduct a precision robustness analysis under different numerical settings. As shown in Table 4, running the perplexity experiments under different numerical precisions gives almost the same results, indicating that these neurons are not numerical outliers but functionally critical components (the detailed analysis can be found in Appendix A.10).
Table 4: Precision robustness test on representative models using 1,000 WikiText-103 samples. Left to right: each column shows the model’s perplexity before and after masking critical neurons under fp16, fp32, and fp64 precision settings. "Neuron number" is the count of masked critical neurons. All experiments use the same masked neuron set per model.
Model Neuron number fp16 PPL change fp32 PPL change fp64 PPL change
[…]
Cooperative Effect of Critical Neurons. To further verify whether the observed collapse is driven by individual neurons or by their collective interaction, we conduct a lesion analysis over all subsets of the four critical neurons identified in two representative models (Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct).
Table 5 reports the perplexity values before and after masking different neuron combinations. The results clearly show that masking any single neuron or any subset of the critical neurons produces mild degradation, while masking the entire set of four critical neurons leads to catastrophic failure with perplexity increases of up to 5–6 orders of magnitude. This confirms that no single neuron or smaller subset can independently cause collapse. Model failure occurs only when all critical neurons are jointly masked, indicating a cooperative dependency that underlies the observed phase-transition behavior. The result also demonstrates that our method accurately identifies the minimal neuron set.
Table 5: Perplexity results after masking different subsets of the critical neurons in two representative models (Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct). Each column shows the combination of masked neurons and the corresponding perplexity after masking. The baseline perplexity before masking is 17.9204 for Llama-3.2-1B and 13.4696 for Llama-3.2-3B.
Model [0,1,2,3] [1,2,3] [0,1,2] [0,1,3] [0,2,3] [1,3] [0,1] [0,3]
Llama-3.2-1B-Instruct 1585570.7467 22.6918 24.8507 23.8316 23.4580 18.9809 19.5680 19.1943
Llama-3.2-3B-Instruct 386563.3177 30.2199 18.6234 17.7793 15.3875 14.1939 14.1580 14.0209
Model [0,2] [1,2] [2,3] [0] [1] [2] [3]
Llama-3.2-1B-Instruct 19.6826 19.3903 19.0778 18.4835 18.3065 18.4174 18.1181
Llama-3.2-3B-Instruct 13.9827 13.9584 13.8611 13.6737 13.6732 13.5353 13.4421